# Size Winding Mechanism beyond Maximum Chaos

### v1

@@ Abstract
where $J_{j_{1}\cdots j_{q}}$ represents random Gaussian variables with a zero mean and a variance given by $\overline{J_{j_{1}\cdots j_{q}}^{2}}=\frac{(q-1)!\mathcal{J}^{2}}{2qN^{q-1}}$. We specifically focus on the large-$q$ limit by firstly taking $N\rightarrow\infty$ and then $q\rightarrow\infty$. In this scenario, both $h^{R}$ and $f^{A}$ have been computed in closed-form in Ref. [44], which leads to
@@
Here, $y=\lambda_{0}^{-1}\left[\cos\Big(\frac{\pi v}{2}\Big{missing})(1-2s)^{-\frac{1}{2\Delta}}-1\right]$, which indicates we have non-vanishing (winding) size distribution only for $s\in[\frac{1-\cos^{2\Delta}(\pi v/2)}{2},\frac{1}{2}]$. We have introduced $\Delta=1/q$ and $\varkappa=2\pi v/\beta$ with $v=\frac{\beta\mathcal{J}}{\pi}\cos\frac{\pi v}{2}$. Interestingly, $\mathcal{P}(s,t)=|\mathcal{Q}(s,t)|$ is satisfied for arbitrary time. Nevertheless, the phase of $\mathcal{Q}(s,t)$ is linear in $y$ instead of $s$. Therefore, the large-$q$ SYK model exhibits perfect size winding only in the early-time regime $t\ll t_{\text{sc}}$ where typical operator has $s\ll 1$ and we can take the approximation $y\approx\lambda_{0}^{-1}\left[\cos\Big(\frac{\pi v}{2}\Big{missing})(1+qs)-1\right]$. This is demonstrated by a numerical plot in FIG. 3. We can further compute the slope of the winding phase as
@@
Long-time limit & finite $N$.– Now, we turn our attention to the long-time limit with $t-t_{\text{sc}}\gg\varkappa^{-1}$. The initial result Eq. (12) naively suggests that the variance of the winding size distribution decays exponentially over time, eventually becoming smaller than the finite-$N$ broadening caused by a finite $\sigma$ in Eq. (8). Furthermore, neglecting this finite-$N$ broadening leads to the divergence of $\text{arg}~{}Q$ near $s=1/2$, as illustrated in FIG 3. Consequently, to obtain the correct result, it is imperative to retain a finite $\sigma$ in Eq. (8). This finite-$N$ correction is particularly crucial in the NISQ era, where quantum teleportation is performed on systems with a small number of qubits [17]. Additionally, it is essential for comparing theoretical predictions with numerics using exact diagonalization (ED), which is provided in the supplementary material [18]. We numerically plot the size and winding size distributions with finite-$N$ corrections, as illustrated in FIG. 4. Firstly, we observe that $\mathcal{P}(s,t)$ and $|\mathcal{Q}(s,t)|$ are not exactly equal, although they qualitatively agree with each other, as seen in FIG. 4(a). $|\mathcal{Q}(n,t)|$ is smaller than $\mathcal{P}(n,t)$, indicating that the phase cancellation within fixed operator length sector appears as a leading finite-$N$ correction. Secondly, due to the finite-$N$ boardening, the peaks alway have finite widths and distribution functions are non-vanishing for the entire region of $s\in[0,1]$. Thirdly, the behavior of the winding phase with finite-$N$ correction undergoes significant changes, showing near-perfect size winding for arbitrary $s$ in the long-time limit, which was absent in the large-$N$ result. We can further expand Eq. (8) with $\lambda_{0}\rightarrow\infty$ to obtain the asymptotic form of the winding phase. For the large-$q$ SYK model, this gives[18]
@@
where $\psi^{0}(2\Delta)$ is the polyGamma function and $\gamma$ is the Euler constant. We find that this approximate formula matches well with Fig. 4(b) and confirms the linear behavior in the $s\sim\mathcal{O}(1)$ region. We have also employed the ED method to calculate the winding phase in the $N=18$ system, which also demonstrates linear behavior in the $s\sim\mathcal{O}(1)$ region [18]. We thereby conclude that the $s\sim\mathcal{O}(1)$ linear winding phase receives important finite-size effects in small-size systems, indicating that teleportation operates via a completely different mechanism compared to the large-$N$ case.
@@
Discussions.– In this study, we examine the distribution of winding sizes as a detailed probe into information scrambling. Our results reveal that the winding phase is inherently linked to the pure imaginary time shift of an operator and its corresponding Keldysh contour. This connection results in a universal phase factor for scramblon propagators. The results show that the large-$q$ SYK model exhibits perfect size winding in the large-$N$ limit during the early-time regime, where typical operators possess a size $s\ll 1$. We further examine the effects of finite-$N$ corrections, which prove to be crucial in the long-time limit. This analysis uncovers a linear winding phase at $s\sim\mathcal{O}(1)$ in small-size systems, aligning with numerical simulations conducted through ED. Several additional remarks are pertinent. Firstly, although our focus is on chaotic quantum systems, recent studies have demonstrated that integrable systems, such as the commuting SYK model [17, 46], can also exhibit size winding in specific parameter regimes. Consequently, it becomes intriguing to inquire about the necessary conditions for the size winding mechanism. Secondly, while the holographic picture or maximal chaos are indispensable for the size winding mechanism, they can still play a crucial role in other facets of the wormhole teleportation protocol, such as the causal time-ordering of teleported signals and the Shapiro time delay. Thirdly, our results emphasize the pervasive influence of finite-$N$ effects in recent research, particularly when utilizing numerical tools or conducting experiments on NISQ quantum devices [17]. Advanced techniques are necessary to accurately observe and interpret theoretical results obtained in the large-$N$ limit. Acknowledgement. We thank Ping Gao, Xiao-Liang Qi, Jinzhao Wang and Ying Zhao for helpful discussions. The project is supported by NSFC under Grant No. 12374477.
Supplemental Material: ”Universal Size Winding Mechanism without Maximum Chaos”
@@ Appendix A Wormhole teleportation and size winding
$P(n,t)=\sum_{j_{1}<j_{2}<\cdots<j_{n}}\left|c_{j_{1}j_{2}\cdots j_{n}}(t)\right|^{2},\ \ \ \ \ \ \ \ \ \ Q(n,t)=\sum_{j_{1}<j_{2}<\cdots<j_{n}}\left(c_{j_{1}j_{2}\cdots j_{n}}(t)\right)^{2}.$ (S1)
We focus on the time-dependent amplitudes $c_{j_{1}j_{2}\cdots j_{n}}(t)$, which become significant when the composite operator $\hat{O}$ incorporates a Heisenberg-evolved operator $\hat{O}(t)$. Typically, $\hat{O}$ is non-Hermitian, resulting in complex amplitudes that can be expressed as $c{j_{1}j_{2}\cdots j_{n}}(t)=|c_{j_{1}j_{2}\cdots j_{n}}(t)|\exp(i\theta_{j_{1}j_{2}\cdots j_{n}}(t))$. This complexity introduces a size-dependent phase in the winding size distribution, which we represent as $\operatorname{arg}Q(n,t)\equiv\theta(n,t)$. Previous studies have established that, ideally, $|Q(n,t)|$ is equivalent to $P(n,t)$, and $\theta(n,t)$ exhibits linear scaling with $n$. The linear scaling is indicative of optimal size winding, a crucial condition for achieving an optimal wormhole teleportation signal [17, 13, 15, 16].
@@
We first note that in Eq. (S2), the term $e^{-igV}$ does not form an out-of-time-order correlator (OTOC) when combined with $\chi_{k}^{R}(t)\chi_{k}^{L}(-t).$ For sufficiently long time $t$, it can be approximately factored out. Therefore, the correlator can be reformulated as follows:
$F(t)=e^{-{i\mkern 1.0mu}g\langle V\rangle}\bra{\text{EPR}}\rho_{\beta,R}^{1/2}\chi_{k}^{R}(t)e^{igV}\rho_{\beta,R}^{1/2}\chi_{k}^{R}(t)\ket{\text{EPR}}.$ (S3)
The expectation of the coupling is defined under the TFD state. Secondly, by specifically choosing the coupling $V=\sum_{j}i\chi_{j}^{L}\chi_{j}^{R},$ we can substantially simplify the expression of the correlator in Eq. (S2). Up to a constant, the coupling operator $V$ matches the size operator, as discussed in the main text following Ref. [47]. Expanding $\rho_{\beta,R}^{1/2}\chi_{k}^{R}(t)$ in terms of Majorana fermion basis $\rho_{\beta,R}^{1/2}\chi_{k}^{R}(t)=\sum_{n}\sum_{j_{1}<j_{2}<\cdots<j_{n}}c_{j_{1}j_{2}\cdots j_{n}}{i\mkern 1.0mu}^{\lfloor n/2\rfloor}\chi_{j_{1}}^{R}\chi_{j_{2}}^{R}\cdots\chi_{j_{n}}^{R}$, the correlator can be related to the winding size distribution.
$F(t)=\sum_{n=0}^{N}e^{-{i\mkern 1.0mu}g\langle V\rangle}e^{{i\mkern 1.0mu}g(2n-N)}Q(n,t),$ (S4)
@@
To attain the maximum value of $F(t)$, two specific conditions are essential. The first condition is that the magnitude of $Q(n,t)$, denoted as $|Q(n,t)|$, should approximate $P(n,t)$, thereby placing constraints on the amplitude. The second condition entails setting $\arg Q(n,t)$ to $2i\alpha n$. This is intended to counterbalance the phase induced by the coupling, represented as $2ign$. Implementing the relation $\alpha+g=0$ and adhering to the normalization condition for the operator, which states that $\sum_{n}P(n,t)=1$, we find that the two-sided correlator attains its peak value, $\max F(t)=e^{-ig(\langle V\rangle-N)}$. This demonstrates an optimal teleportation signal when the system exhibits perfect size winding [17, 13, 15, 16].
Appendix B The derivation of generating function with finite-$N$ correction
As explained in the main text, the generating functions can be expressed as
$\mathcal{S}_{\mathcal{P}/\mathcal{Q}}=e^{-\frac{\nu}{2}+\frac{\nu^{2}}{8N}}\left<\text{T}_{c}\chi_{k}(T_{1})\chi_{k}(T_{2})\prod_{j=1}^{N}{\left(1+\tanh(\frac{\nu}{2N})\chi_{j}(T_{3})\chi_{j}(T_{4})\right)}\right>.$ (S5)
In this section, we present the details of the scramblon calculation. We expand the product of Majorana operators to the Majorana string basis, which leads to
$\begin{split}\prod_{j=1}^{N}{\left(1+\tanh(\frac{\nu}{2N})\chi_{j}(T_{3})\chi_{j}(T_{4})\right)}&=\sum_{n=1}^{N}\sum_{|\bm{J}|=n}\left((-1)^{\frac{(n-1)n}{2}}\chi_{\bm{J}}(T_{3})\chi_{\bm{J}}(T_{4})\tanh(\frac{\nu}{2N})^{n}\right)\\ \end{split}.$ (S6)
Our approach to compute the generating function is analogous to the calculation of the OTOC as a summation of scramblon modes, detailed in Eq. (6) of the main text.
$\begin{split}\mathcal{S}_{\mathcal{P}/\mathcal{Q}}&=e^{-\frac{\nu}{2}+\frac{\nu^{2}}{8N}}\sum_{n=0}^{N}\sum_{|\bm{J}|=n}\sum_{m_{1},\cdots,m_{n}}\Upsilon^{{\mathrm{R}},\sum_{l=1}^{n}m_{l}}(T_{12})\left(\frac{(-\lambda)^{m_{1}}}{m_{1}!}\Upsilon^{{\mathrm{A}},m_{1}}(T_{34})\cdots\frac{(-\lambda)^{m_{n}}}{m_{n}!}\Upsilon^{{\mathrm{A}},m_{n}}(T_{34})\right)\tanh(\frac{\nu}{2N})^{n}.\end{split}$ (S7)
It is important to note that the retarded vertex can be expressed through the function $h^{R}$.
$\Upsilon^{{\mathrm{R}},\sum_{l}m_{l}}(T_{12})=\int_{0}^{\infty}\mathop{}\!\mathrm{d}yy^{\sum_{l}m_{l}}h^{R}(y,T_{12})$ (S8)
and we can combine the introduced variable $y$ to scramblon propagator $\lambda$ and the advanced vertex to get another function $f^{A}$
$\sum_{l}\frac{(-\lambda y)^{m_{l}}}{m_{l}!}\Upsilon^{{\mathrm{A}},m_{l}}(T_{34})=\sum_{l}\int_{0}^{\infty}\mathop{}\!\mathrm{d}y_{l}\frac{\left(-\lambda yy_{l}\right)^{m_{l}}}{m_{l}!}h^{A}(y,T_{34})=\int_{0}^{\infty}\mathop{}\!\mathrm{d}y_{l}e^{-\lambda yy_{l}}h^{A}(y_{l},T_{34})=f^{A}(\lambda y,T_{34})$ (S9)
By transforming the vertex function to $h^{R}$ and $f^{A}$ function, finally we arrive at
$\begin{split}\mathcal{S}_{\mathcal{P}/\mathcal{Q}}&=e^{-\frac{\nu}{2}+\frac{\nu^{2}}{8N}}\int_{0}^{\infty}\mathop{}\!\mathrm{d}yh^{R}(y,T_{12})\sum_{n=0}^{N}\sum_{|\bm{J}|=n}\left(f^{A}(\lambda y,T_{34})\right)^{n}\tanh(\frac{\nu}{2N})^{n}\\ &=e^{-\frac{\nu}{2}+\frac{\nu^{2}}{8N}}\int_{0}^{\infty}\mathop{}\!\mathrm{d}yh^{R}(y,T_{12})\left(1+f^{A}(\lambda y,T_{34})\tanh(\frac{\nu}{2N})\right)^{N}\end{split}$ (S10)
Eq.(S10) represents the precise generating function within the scramblon theory framework, forming the foundational equation for our study. In this equation, the term $-i\chi_{j}^{L}\chi_{j}^{R}$ is substituted with $f^{A}$ as shown in Eq.(S5). This substitution reiterates the physical significance of $f^{A}$, which characterizes the fermion two-point function influenced by scramblon perturbation of strength $x$.
Treating $\frac{\nu}{2N}$ as a small parameter, we expand the expression to the second order of $\left(\frac{\nu}{2N}\right)^{2}$ in the exponential term. However, the final expansion is taken only to the leading order of $\mathcal{O}(\frac{1}{N})$, resulting in the following expression.
$\begin{split}\mathcal{S}(\nu,t)_{\mathcal{P}/\mathcal{Q}(s,t)}&=\int_{0}^{+\infty}\mathop{}\!\mathrm{d}y\,h^{\mathrm{R}}(y,T_{12})e^{\frac{\nu^{2}}{8N}\left(1-f^{A}(\lambda y,T_{34})^{2}\right)-\frac{\nu}{2}\left(1-f^{A}(\lambda y,T_{34})\right)}\\ \end{split}$ (S11)
By the inverse Laplace transform, we can obtain the size distribution or winding size distribution
$\mathcal{P}/\mathcal{Q}(s,t)=\int_{0}^{+\infty}\mathop{}\!\mathrm{d}y\,h^{\mathrm{R}}(y,T_{12})\frac{1}{\sqrt{2\pi\sigma^{2}}}e^{-\frac{1}{2\sigma^{2}}\left(s-\frac{1-f^{A}(\lambda y,T_{34})}{2}\right)^{2}},\ \ \ \ \ \ \ \ \ \ \ \text{with}\ \sigma^{2}=\frac{1-f^{A}(\lambda y,T_{34})^{2}}{4N}.$ (S12)
In the large-$N$ limit, the finite width of the Gaussian function narrows to zero, leading to its replacement with a $\delta$ function, centered at the same position.
Appendix C SYK model and scramblon calculation in large-$N$ and finite-$N$ limit
In this section, we present the explicit formula for the distribution functions in the SYK model. In the large-$q$ SYK model, $h^{R},f^{A}$ can be explicitly evaluated in terms of SYK parameter [19]
$\begin{split}h^{R}(y,T_{12})&=\frac{y^{2\Delta-1}\cos^{2\Delta}\left(\frac{\pi v}{2}\right)\exp\left(-y\cos\left(\pi v\left(\frac{1}{2}-\frac{{i\mkern 1.0mu}T_{12}}{\beta}\right)\right)\right)}{\Gamma(2\Delta)}\\ f^{A}(\lambda y,T_{34})&=\cos^{2\Delta}\left(\frac{\pi v}{2}\right)\left(\cos\left(\pi v\left(\frac{1}{2}-\frac{{i\mkern 1.0mu}T_{34}}{\beta}\right)\right)+\lambda y\right)^{-2\Delta},\\ \end{split}$ (S13)
Notice that here $\lambda$ also depends on time argument $T_{1}\cdots T_{4}$ in the growing Lyapunov exponent
$\lambda=C^{-1}e^{{i\mkern 1.0mu}\varkappa\frac{\beta}{2}+\varkappa\frac{T_{1}+T_{2}-T_{3}-T_{4}}{2}}$ (S14)
Large-$N$
By using Eq. (9) in the main text, it leads to the final result of the operator distribution function
$\begin{split}\mathcal{P}(s,t)&=\frac{2y_{0}^{2\Delta-1}\exp\left(-\cos\left(\frac{\pi v}{2}\right)y_{0}\right)(1-2s)^{-\frac{2\Delta+1}{2\Delta}}}{\lambda_{0}\Gamma(2\Delta+1)}\\ \mathcal{Q}(s,t)&=\frac{2y_{0}^{2\Delta-1}\exp\left(-e^{-\frac{{i\mkern 1.0mu}\pi v}{2}}y_{0}\right)(1-2s)^{-\frac{2\Delta+1}{2\Delta}}}{\lambda_{0}\Gamma(2\Delta+1)}e^{-{i\mkern 1.0mu}\Delta\pi v}\\ &\text{with}\ y_{0}=\lambda_{0}^{-1}\left[\cos\Big(\frac{\pi v}{2}\Big{missing})(1-2s)^{-\frac{1}{2\Delta}}-1\right],\ \lambda_{0}=C^{-1}e^{\varkappa t}.\\ \end{split}$ (S15)
We find that $\mathcal{P}(s,t)=|\mathcal{Q}(s,t)|$ exactly, as stated in the main text.
Finite-$N$
In the finite-$N$ case, an analytical formula generally does not exist. Nonetheless, the linear winding phase behavior becomes evident in the long-time limit. In this regime, the function $f^{A}(\lambda y,T_{34})$ approaches zero, as indicated by Eq. (S13). This observation justifies a series expansion in terms of $f^{A}(\lambda y,T_{34})$. At the leading order, the expansion yields
$\mathcal{P}/\mathcal{Q}(s,t)=\int_{0}^{+\infty}\mathop{}\!\mathrm{d}y\,h^{\mathrm{R}}(y,T_{12})\sqrt{\frac{2N}{\pi}}e^{-2N\left(s-\frac{1}{2}\right)^{2}}\left(1+Nf^{A}(\lambda y,T_{34})(1-2s)\right)+O\left(f^{A}(\lambda y,T_{34})^{2}\right)$ (S16)
For a detailed examination of the winding phase behavior, we set $T_{12}=-{i\mkern 1.0mu}\beta/2,T_{34}=-{i\mkern 1.0mu}\beta/2$ and $\lambda=C^{-1}e^{\varkappa t+{i\mkern 1.0mu}\varkappa\frac{\beta}{4}}=C^{-1}e^{2\pi v\frac{t}{\beta}+\frac{{i\mkern 1.0mu}\pi v}{2}}$, as prescribed by the $S_{\mathcal{Q}}$ contour.
$\mathcal{Q}(s,t)=\sqrt{\frac{2N}{\pi}}e^{-\frac{1}{2}N(1-2s)^{2}}\cos^{2\Delta}\left(\frac{\pi v}{2}\right)\left(1+N(1-2s)\mathcal{U}\left(2\Delta,1,\frac{1}{\lambda}\right)\left(\frac{\cos\left(\frac{\pi v}{2}\right)}{\lambda}\right)^{2\Delta}\right)$ (S17)
[…]
To estimate the winding phase, we approximate Eq. (S17) using $1+x\approx e^{x}$ for small $x$. Consequently, the winding phase is deduced by taking the imaginary part of the exponent:
$\begin{split}\arg\mathcal{Q}(s,t)&=\operatorname{Im}N(1-2s)\mathcal{U}\left(2\Delta,1,\frac{1}{\lambda}\right)\left(\frac{\cos\left(\frac{\pi v}{2}\right)}{\lambda}\right)^{2\Delta}\\ &\approx(2s-1)\frac{N}{\Gamma(2\Delta)}\left(\cos\left(\frac{\pi v}{2}\right)e^{-\frac{2\pi tv}{\beta}}\right)^{2\Delta}\left(\sin(\pi\Delta v)\left(\frac{2\pi tv}{\beta}-\psi^{(0)}(2\Delta)-2\gamma\right)-\frac{1}{2}\pi v\cos(\pi\Delta v)\right).\\ \end{split}$ (S18)
This analysis confirms that, even in a finite-$N$ system, the late-time winding phase is proportional to $12s-$ 1, exhibiting an exponential decay in time. This implies the persistence of the winding mechanism at late times due to a crucial finite-$N$ correction.
Appendix D Exact Diagonalization Numerics

### v2

@@ 3 Size winding at large-$N$
where $J_{j_{1}\cdots j_{q}}$ represents random Gaussian variables with a zero mean and a variance given by $\overline{J_{j_{1}\cdots j_{q}}^{2}}=\frac{(q-1)!\mathcal{J}^{2}}{2qN^{q-1}}$. We specifically focus on the large-$q$ limit by firstly taking $N\rightarrow\infty$ and then $q\rightarrow\infty$. In this scenario, both $h^{R}$ and $f^{A}$ have been computed in closed-form in Ref. [43]:
$\begin{split}h^{R}(y,T_{12})&=\frac{y^{2\Delta-1}\cos^{2\Delta}\left(\frac{\pi v}{2}\right)\exp\left(-y\cos\left(\pi v\left(\frac{1}{2}-\frac{{i\mkern 1.0mu}T_{12}}{\beta}\right)\right)\right)}{\Gamma(2\Delta)},\\ f^{A}(\lambda y,T_{34})&=\cos^{2\Delta}\left(\frac{\pi v}{2}\right)\left(\cos\left(\pi v\left(\frac{1}{2}-\frac{{i\mkern 1.0mu}T_{34}}{\beta}\right)\right)+\lambda y\right)^{-2\Delta}.\\ \end{split}$ (17)
We have introduced $\Delta=1/q$ and $\varkappa=2\pi v/\beta$ with $v=\frac{\beta\mathcal{J}}{\pi}\cos\frac{\pi v}{2}$. This leads to
@@
Here, $y=\lambda_{0}^{-1}\left[\cos\Big(\frac{\pi v}{2}\Big{missing})(1-2s)^{-\frac{1}{2\Delta}}-1\right]$, which indicates we have non-vanishing (winding) size distribution only for $s\in[\frac{1-\cos^{2\Delta}(\pi v/2)}{2},\frac{1}{2}]$. Interestingly, $\mathcal{P}(s,t)=|\mathcal{Q}(s,t)|$ is satisfied for arbitrary time. Nevertheless, the phase of $\mathcal{Q}(s,t)$ is linear in $y$ instead of $s$. Therefore, the large-$q$ SYK model exhibits perfect size winding only in the early-time regime $t\ll t_{\text{sc}}$ where typical operator has $s\ll 1$ and we can take the approximation $y\approx\lambda_{0}^{-1}\left[\cos\Big(\frac{\pi v}{2}\Big{missing})(1+qs)-1\right]$. This is demonstrated by a numerical plot in FIG. 3. We can further compute the slope of the winding phase as
@@
4 Long-time limit & finite $N$
Now, we turn our attention to the long-time limit with $t-t_{\text{sc}}\gg\varkappa^{-1}$. The initial result Eq. (18) naively suggests that the variance of the winding size distribution decays exponentially over time, eventually becoming smaller than the finite-$N$ broadening caused by a finite $\sigma$ in Eq. (13). Furthermore, neglecting this finite-$N$ broadening leads to the divergence of $\text{arg}~{}Q$ near $s=1/2$, as illustrated in FIG. 3. Consequently, to obtain the correct result, it is imperative to retain a finite $\sigma$ in Eq. (13). This finite-$N$ correction is particularly crucial in the NISQ era, where quantum teleportation is performed on systems with a small number of qubits [17]. Additionally, it is essential for comparing theoretical predictions with numerics using exact diagonalization (ED), which is provided in the Appendix B.
We numerically plot the size and winding size distributions with finite-$N$ corrections, as illustrated in FIG. 4. Firstly, we observe that $\mathcal{P}(s,t)$ and $|\mathcal{Q}(s,t)|$ are not exactly equal, although they qualitatively agree with each other, as seen in FIG. 4(a). $|\mathcal{Q}(n,t)|$ is smaller than $\mathcal{P}(n,t)$, indicating that the phase cancellation within fixed operator length sector appears as a leading finite-$N$ correction. Secondly, due to the finite-$N$ boardening, the peaks alway have finite widths and distribution functions are non-vanishing for the entire region of $s\in[0,1]$. Thirdly, the behavior of the winding phase with finite-$N$ correction undergoes significant changes, showing near-perfect size winding for arbitrary $s$ in the long-time limit, which was absent in the large-$N$ result. We can further expand Eq. (13) with $\lambda_{0}\rightarrow\infty$ to obtain the asymptotic form of the winding phase. For the large-$q$ SYK model, this gives
@@ 4 Long-time limit & finite $N$
where $\psi^{0}(2\Delta)$ is the polyGamma function and $\gamma$ is the Euler constant. We find that this approximate formula matches well with Fig. 4(b) and confirms the linear behavior in the $s\sim\mathcal{O}(1)$ region. In Appendix B, we have also employed the ED method to calculate the winding phase in the $N=18$ system, which also demonstrates linear behavior in the $s\sim\mathcal{O}(1)$ region. We thereby conclude that the $s\sim\mathcal{O}(1)$ linear winding phase receives important finite-size effects in small-size systems, indicating that teleportation operates via a completely different mechanism compared to the large-$N$ case.
@@
5 Discussion
In this study, we examine the distribution of winding sizes as a detailed probe into information scrambling. Our results reveal that the winding phase is inherently linked to the pure imaginary time shift of an operator and its corresponding Keldysh contour. This connection results in a universal phase factor for scramblon propagators. The results show that the large-$q$ SYK model exhibits perfect size winding in the large-$N$ limit during the early-time regime, where typical operators possess a size $s\ll 1$. We further examine the effects of finite-$N$ corrections, which prove to be crucial in the long-time limit. This analysis uncovers a linear winding phase at $s\sim\mathcal{O}(1)$ in small-size systems, aligning with numerical simulations conducted through ED.
Several additional remarks are pertinent. Firstly, although our focus is on chaotic quantum systems, recent studies have demonstrated that integrable systems, such as the commuting SYK model [17, 45], can also exhibit size winding in specific parameter regimes. Consequently, it becomes intriguing to inquire about the necessary conditions for the size winding mechanism. Secondly, while the holographic picture or maximal chaos are indispensable for the size winding mechanism, they can still play a crucial role in other facets of the wormhole teleportation protocol, such as the causal time-ordering of teleported signals and the Shapiro time delay. Thirdly, our results emphasize the pervasive influence of finite-$N$ effects in recent research, particularly when utilizing numerical tools or conducting experiments on NISQ quantum devices [17]. Advanced techniques are necessary to accurately observe and interpret theoretical results obtained in the large-$N$ limit.
Acknowledgment
We thank Ping Gao, Xiao-Liang Qi, Jinzhao Wang and Ying Zhao for helpful discussions. The project is supported by NSFC under Grant No. 12374477 and the National Key R&D Program of China 2023YFA1406702.
@@ Appendix A Wormhole teleportation and size winding
$P(n,t)=\sum_{j_{1}<j_{2}<\cdots<j_{n}}\left|c_{j_{1}j_{2}\cdots j_{n}}(t)\right|^{2},\ \ \ \ \ \ \ \ \ \ Q(n,t)=\sum_{j_{1}<j_{2}<\cdots<j_{n}}\left(c_{j_{1}j_{2}\cdots j_{n}}(t)\right)^{2}.$ (21)
We focus on the time-dependent amplitudes $c_{j_{1}j_{2}\cdots j_{n}}(t)$, which become significant when the composite operator $\hat{O}$ incorporates a Heisenberg-evolved operator $\hat{O}(t)$. Typically, $\hat{O}$ is non-Hermitian, resulting in complex amplitudes that can be expressed as
$c_{j_{1}j_{2}\cdots j_{n}}(t)=|c_{j_{1}j_{2}\cdots j_{n}}(t)|\exp(i\theta_{j_{1}j_{2}\cdots j_{n}}(t)).$
This complexity introduces a size-dependent phase in the winding size distribution, which we represent as $\operatorname{arg}Q(n,t)\equiv\theta(n,t)$. Previous studies have established that, ideally, $|Q(n,t)|$ is equivalent to $P(n,t)$, and $\theta(n,t)$ exhibits linear scaling with $n$. The linear scaling is indicative of optimal size winding, a crucial condition for achieving an optimal wormhole teleportation signal [13, 15, 16, 17].
@@
We first note that in Eq. (22), the term $e^{-igV}$ does not form an out-of-time-order correlator (OTOC) when combined with $\chi_{k}^{R}(t)\chi_{k}^{L}(-t).$ For sufficiently long time $t$, it can be approximately factored out. Therefore, the correlator can be reformulated as follows:
$F(t)=e^{-{i\mkern 1.0mu}g\langle V\rangle}\bra{\text{EPR}}\rho_{\beta,R}^{1/2}\chi_{k}^{R}(t)e^{igV}\rho_{\beta,R}^{1/2}\chi_{k}^{R}(t)\ket{\text{EPR}}.$ (23)
The expectation of the coupling is defined under the TFD state. Secondly, by specifically choosing the coupling $V=\sum_{j}i\chi_{j}^{L}\chi_{j}^{R},$ we can substantially simplify the expression of the correlator in Eq. (22). Up to a constant, the coupling operator $V$ matches the size operator, as discussed in the main text following Ref. [46]. Expanding $\rho_{\beta,R}^{1/2}\chi_{k}^{R}(t)$ in terms of Majorana fermion basis $\rho_{\beta,R}^{1/2}\chi_{k}^{R}(t)=\sum_{n}\sum_{j_{1}<j_{2}<\cdots<j_{n}}c_{j_{1}j_{2}\cdots j_{n}}{i\mkern 1.0mu}^{\lfloor n/2\rfloor}\chi_{j_{1}}^{R}\chi_{j_{2}}^{R}\cdots\chi_{j_{n}}^{R}$, the correlator can be related to the winding size distribution.
$F(t)=\sum_{n=0}^{N}e^{-{i\mkern 1.0mu}g\langle V\rangle}e^{{i\mkern 1.0mu}g(2n-N)}Q(n,t),$ (24)
@@
To attain the Maximal value of $F(t)$, two specific conditions are essential. The first condition is that the magnitude of $Q(n,t)$, denoted as $|Q(n,t)|$, should approximate $P(n,t)$, thereby placing constraints on the amplitude. The second condition entails setting $\arg Q(n,t)$ to $2i\alpha n$. This is intended to counterbalance the phase induced by the coupling, represented as $2ign$. Implementing the relation $\alpha+g=0$ and adhering to the normalization condition for the operator, which states that $\sum_{n}P(n,t)=1$, we find that the two-sided correlator attains its peak value, $\max F(t)=e^{-ig(\langle V\rangle-N)}$. This demonstrates an optimal teleportation signal when the system exhibits perfect size winding [13, 15, 16, 17].
Appendix B Exact Diagonalization Numerics